OpenStudy (anonymous):

Tom has a collection of 30 CDs and Nita has a collection of 18 CDs. Tom is adding 1 CD a month to his collection while Nita is adding 5 CDs a month to her collection. Find the number of months after which they will have the same number of CDs. A. 1 month B. 3 months C. 2 months D. 33 months

OpenStudy (anonymous):

\[30+t=18+5t\\12=4t\\t=3\]